i know this is basic math but i am lost the teacher wants us to solve for the unknown. i do not even know where to begin. can you help me please?
s/3.0 = 1.0/0.2
i am clueless

Answer by Annabelle1(69)

s/3.0 = 1.0/0.2
to get s as the subject of the formula we need to multiply both sides of the equation by 3.0.
(s/3.0)*3.0=(1.0/0.2)*3.0
s= 5*3.0 (1.0/0.2=5)
s=15